OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10) and (-4, 0)?

OpenStudy (ganpat):

slope m = (y2 - y1) / (x2 - x1) where x1, y1 & x2, y2 are given points...

OpenStudy (ganpat):

m = (0- (-10)) / (-4 - 0) m = 10/-4 = -5/2...

OpenStudy (anonymous):

\[slope=DeltaY/DeltaX ;then: slope=(Y2-Y1)/(X2-X1)\]